# Quantifying the dynamic structural resilience of international staple food trade networks: An entropy-based approach

### v2

@@ Abstract
Establishing a resilient food trade system is an international consensus on safeguarding food security amid growing disruptions. We employ an entropy-based approach to quantify the dynamic structural resilience of international trade networks for maize, rice, soybean, and wheat from 1986 to 2022. Using index decomposition analysis, we also investigate the relative contributions of several internal components to changes in resilience. Within this framework, despite heterogeneity across different food commodities, we find that current trade networks are relatively redundant, with improvements in efficiency being the dominant driver of changes in resilience. In addition, we reveal a historically pronounced impact of flow concentrations on resilience, while trade interactions have become increasingly important in recent years. Furthermore, following the leave-one-out approach, we identify critical economies and trade relationships that disproportionately affect the overall resilience, many of which are overlooked by conventional volume-based or centrality-based metrics. Moreover, we highlight that the overconcentration of flows along core trade relationships may undermine the overall efficiency and resilience, whereas peripheral trade networks play an essential role in sustaining resilience, underscoring the importance of promoting more equitable trade relations. These findings not only provide new insights for assessing the resilience of international food trade systems but also propose directions for making them more resilient.
@@ 1 Introduction
Resilience is a fundamental property of systems (Liu et al., 2022), yet it remains a contested construct, with ongoing and often discipline-specific debates among various schools of thought, including those in engineering, psychology, management, and economics (Youssef and Luthans, 2007; Francis and Bekera, 2014; Brunnermeier, 2024; Padovano and Ivanov, 2025; Gunasekaran et al., 2015; Ivanov, 2024). Its concept originates in ecology, where Holling (1973) defines it as a system’s capacity to withstand disturbances without collapsing or transitioning to a fundamentally altered state, and to recover its essential functions afterward. Building on this concept, scholars initially define food system resilience as the ability to maintain a stable food supply—both in terms of quantity and diversity—while withstanding disruptions (Tendall et al., 2015). As the food trade system, accounting for approximately 20% of the food consumed globally, constitutes a critical subsystem of the broader food system (D’Odorico et al., 2014), its resilience is often conceptualized in a similar manner (Mena et al., 2022; Jafari et al., 2024).
@@
While efforts to accurately define the resilience of the food trade system continue, scholars also initiate discussions on how it can be effectively measured, yet no consensus emerges (Pimm et al., 2019). Most resilience measures are based on the concept of the resilience triangle (see Fig. 1 in Bruneau et al. (2003)), which defines resilience as the level of system performance retained during a disturbance period (Cheng et al., 2022). For example, studies in engineering often assess resilience by measuring the cumulative loss in infrastructure quality over time (Fang et al., 2022; Chen et al., 2024; Sahebjamnia et al., 2015). In economics, resilience is often understood as an economic system’s dynamic response to external shocks, and can be thus measured through response functions (Klimek et al., 2019; Brunnermeier, 2024). In ecology, resilience is typically quantified in terms of recovery rate or the time required to adapt to changing disturbance regimes (Seidl et al., 2016). Accordingly, Wei et al. (2025) extend this concept to “V-shaped curve” and assess resilience of food trade networks from a port perspective. These studies emphasize the system’s speed of recovery and the magnitude of loss, offering valuable insights into resilience measurement. However, these approaches may be inadequate, as they focus primarily on system performance following a disturbance while neglecting that in the absence of shocks, which may not completely reflect the fact that resilience is a fundamental system property (Liu et al., 2022). Meanwhile, determining the duration of shocks and the appropriate recovery level remains challenging and unrealistic (Hosseini et al., 2016).
@@
Most real-world systems are structured as networks with complex topologies (Barabási and Albert, 1999). Drawing on approaches from network science, a substantial body of literature assesses system resilience by analyzing topological metrics of the network. For example, resilience can be assessed using the inverse of the harmonic mean of the shortest path lengths between nodes, referred to as global efficiency (which differs from the notion of efficiency used in our study) (Latora and Marchiori, 2001; Osei-Asamoah and Lownes, 2014; Testa et al., 2015; Xie et al., 2021; Wei et al., 2022; Chen et al., 2024), which is also applied to assess resilience in the global agri-food wheat trade system (Fair et al., 2017). Moreover, Dixit et al. (2020) measure supply chain resilience by integrating four topological properties, including density, centrality, size, and connectivity, which is also applied to assess resilience in crude oil and gas trade networks (Mou et al., 2020b; Sun et al., 2023). Although topological properties primarily capture the structural aspects of a networked system, they overlook its dynamic characteristics (Qi and Mei, 2024) and do not adequately reflect the actual dynamic evolution of system resilience, particularly in food trade system that is inherently directed and weighted. Notably, within a framework proposed by Gao et al. (2016), existing literature can infer whether a system is resilient based on dynamic equations (Xu et al., 2019; Zhao et al., 2024; Tu et al., 2019), which is also applied to food trade systems (Wei et al., 2025). However, this resilience inference has to rely on the strong assumptions that the degrees of interconnected nodes are largely independent (Liu et al., 2024a). These assumptions contrast with the high assortativity observed in international trade networks (Mou et al., 2020a), which may limit the applicability of such framework in this context.
Information theory is introduced to fill the gaps in measuring resilience of trade systems. Although resilience encompasses several principles such as adaptability and cohesion (Bhamra et al., 2011), diversification is widely regarded as the most fundamental, as it manifests at multiple levels and across different components of the food system (Hertel et al., 2021). Higher diversification implies a broader distribution of probabilities, indicating that the food system is less likely to collapse in the face of disturbances (Nicholson et al., 2021). In this context, entropy, an information-theoretic measure of uncertainty and diversity, provides a more effective representation of system resilience, particularly for systems characterized by directed and weighted flows. Among related studies, entropy is first applied to assess the stability of food webs in ecosystems (Rutledge et al., 1976; Ulanowicz, 1979). Specifically, based on the entropy-based approach, the trade-off between two underlying system properties—ascendency and reserve—is proposed to measure resilience of ecosystems (Ulanowicz, 2009; Ulanowicz et al., 2009). Kharrazi et al. (2017) argue that flows of energy in food webs are inherently exchanges of resources and subsequently assess resilience of global commodity trade systems by redefining the above two properties as efficiency and redundancy. Building on this line of research, scholars recently begin to apply this approach to measure resilience in various complex systems, such as phosphorus cycling networks (Liang et al., 2020), coupled nitrigen-phosphorus cycling networks (Luo et al., 2024), waste metals trade systems (Zisopoulos et al., 2025), and power systems (Panyam et al., 2019), offering interesting insights for our understanding.
Overall, there are substantial academic discussions on the definitions and metrics of resilience, with many of these concepts applied to food systems. However, in our view, research specifically addressing the resilience of food trade systems remains underdeveloped, particularly in the context of increasing geopolitical conflicts and climate risks nowadays. First, most existing studies equate resilience with some unweighted topological properties of the food trade system, which may overlook its dynamic characteristics and the influence of trade volume on overall resilience. Few studies employ information-theoretic entropy to measure the resilience of food trade systems (Li et al., 2024), highlighting a critical gap that warrants further attention. Second, existing studies primarily examine the correlations between the resilience of food trade systems and determinants such as topological properties and external shocks (Li et al., 2024; Ji et al., 2024). However, few provide recommendations for enhancing resilience based on an analysis of its internal components derived through decomposition methods. Such decomposition facilitates a detailed attribution of system resilience to its structural and functional components, thereby revealing its driving mechanisms. Third, although some studies investigate the importance of individual economies within the trading system (Grassia et al., 2022), most rely on indicators such as trade volume and centrality (Gutiérrez-Moya et al., 2021; Zhang and Zhou, 2023), potentially overlooking important nodes characterized by moderate trade volumes and fewer trading partners but playing pivotal structural roles. Moreover, link-level disruptions are more frequent in real-world trade systems, yet existing research predominantly focuses on the significance of nodes, while the importance of trade links remains relatively underexplored (Qian et al., 2017; Guo et al., 2025). Thus, assessing their importance is necessary, especially from a resilience perspective.
@@ 3. RQ3:
The subsequent sections of this paper are structured in the following manner. The methodological framework is presented in Section 2. Section 3 describes the international food trade network and its resilience. Then, Section 4 and Section 5 present and discuss the main findings. Finally, Section 6 serves as the conclusion, which includes the contributions, implications, limitations, and future outlook.
@@ 2 Methodology
2.1 Topological properties of the international food trade system
The international food trade system is usually regarded as a network that exhibits small-world (e.g., high clustering coefficients coupled with short average path lengths) and scale-free (e.g., power-law degree distributions) topological properties in economies’ connectivity (Fagiolo et al., 2009). A food trade network is a weighted digraph $G=\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V}$ represents the vertex set with $N_{\mathcal{V}}$ economies and $\mathcal{E}$ represents the link set with $N_{\mathcal{E}}$ trade relationships. $f_{ij}$ denotes the volume of trade flow that economy $i$ exports to economy $j$. There are several commonly used topological metrics for characterizing network structure, which are briefly described below.
@@ 2.1 Topological properties of the international food trade system
    Modularity represents the effectiveness of partitioning a network into communities, defined as $Q=\frac{1}{2N_{\mathcal{V}}\langle s\rangle}\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\left(f_{ij}-\frac{s_{i}s_{j}}{2N_{\mathcal{V}}\langle s\rangle}\right)\delta(c_{i},c_{j})$, where $\delta(c_{i},c_{j})=1$ if economies $i$ and $j$ belong to the same community and $\delta(c_{i},c_{j})=0$ otherwise. The Louvain algorithm, recognized as one of the fastest and best performing algorithms in comparative analysis (Blondel et al., 2008; Lancichinetti and Fortunato, 2009), is employed for community detection in this paper.
@@ 2.2 Efficiency and redundancy of the international food trade system
In information theory, entropy measures a system’s uncertainty (or diversity). A higher entropy indicates a greater unpredictability and a higher diversity in this system’s states. Accordingly, we can adopt joint entropy to characterize a system’s overall structural properties (Rutledge et al., 1976; Ulanowicz, 1979), which is also called as “capacity” for system development (Ulanowicz and Norden, 1990). Given that joint entropy is equal to the sum of mutual information and conditional entropy, we further introduce two underlying quantities, efficiency and redundancy, which demonstrate opposing properties of a system. Efficiency embodies the flow articulation or constraints within networked configurations, which tends to increase due to preferential interactions between economies (Kharrazi et al., 2017). As economies become more interdependent, trade flows tend to become more concentrated and specialized, thereby enhancing overall system efficiency. In information theory, a higher mutual information indicates a stronger statistical inter-dependence between variables. Let $s=\sum\limits_{i=1}^{N_{\mathcal{V}}}s_{i}^{\mathrm{out}}=\sum\limits_{j=1}^{N_{\mathcal{V}}}s_{j}^{\mathrm{in}}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}f_{ij}$, the efficiency $e$ can be expressed as:
@@
which is also called as the average mutual information (Rutledge et al., 1976; Ulanowicz and Norden, 1990), where
@@
denotes the concentration degree of trade flows from economy $i$ to $j$. $p_{i}^{\mathrm{out}}$ and $p_{j}^{\mathrm{in}}$ respectively denote the concentration degree of trade outflows from economy $i$ to the others:
@@
and that of trade inflows from the others to economy $j$:
@@
Redundancy embodies the diversity of food trade pathways, which is critical for the system’s capacity adapting to changing environmental conditions arising from shocks or disturbances (Luo et al., 2024). A greater number of pathways in food trade allows economies to have greater freedom in the trade market and enhances the system’s robustness in coping with risks. In information theory, conditional entropy measures the uncertainty remaining given that some information is known. A higher conditional entropy indicates a greater availability of alternative pathways in food trade and a consequently higher system redundancy. Hence, the redundancy $r$ can be expressed as
$r&=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{j\mid i}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{i\mid j}\\ &=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{j}^{\mathrm{in}}}\\ &=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{f_{ij}}{s}\ln\frac{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}}{f_{ij}^{2}}\end{aligned}.$ (5)

### v3

@@ Abstract
Establishing a resilient food trade system is an international consensus on safeguarding food security amid growing disruptions. However, a unified resilience framework has yet to be established, leading to the proliferation of diverse measures. Here, we conceptualize resilience as a trade-off between efficiency and redundancy and employ an entropy-based approach to quantify the dynamic structural resilience of international trade networks for maize, rice, soybean, and wheat from 1986 to 2022. Using index decomposition analysis, we also investigate the relative contributions of internal components to resilience dynamics. Within this framework, despite heterogeneity across different food commodities, we find that current trade networks are relatively redundant, with improvements in efficiency being the dominant driver of changes in resilience. In addition, we reveal a historically pronounced impact of flow concentrations on resilience, while trade interactions have become increasingly important in recent years. Following the leave-one-out approach, we furthermore identify critical economies and trade relationships that disproportionately affect the overall resilience, some of which are less well-focused in previous studies. Moreover, we highlight that overconcentration of flows along core trade relationships may undermine both efficiency and resilience, whereas peripheral trade networks may play strategic alternative roles in sustaining resilience, underscoring the importance of concentrating on developing economies and promoting broader trade links. These findings not only provide new insights for assessing the resilience of international food trade systems but also propose directions for strengthening resilience through both regional cooperation and more inclusive trade relations.
@@ 1 Introduction
Resilience is a fundamental property of complex systems (Liu et al., 2022) and its concept originates in ecology, where Holling (1973) defines it as a system’s capacity to withstand disturbances without collapsing or transitioning to a fundamentally altered state, and to recover its essential functions afterward. Building on this concept, various disciplines—including engineering, psychology, management, and economics—have developed discipline-specific definitions of resilience (Youssef and Luthans, 2007; Francis and Bekera, 2014; Brunnermeier, 2024; Padovano and Ivanov, 2025; Gunasekaran et al., 2015; Ivanov, 2024b). However, a unified theory and framework have yet to be established (Liu et al., 2022), which has led to the proliferation of diverse resilience measures (Pimm et al., 2019).
@@
Many scholars emphasize resilience as the capacity of systems to adapt and recover under disruptions (Haimes, 2009), which is commonly assessed using performance-based approaches. These approaches focus primarily on a system’s performance before and after disruptions, regardless of its structural features (Hosseini et al., 2016; Qi and Mei, 2024). Among them, the Resilience Triangle is a widely used framework (Bruneau et al., 2003), which measures resilience as the level of performance retained during a disturbance and visualizes the magnitude, duration, and recovery of performance loss (Cheng et al., 2022). Within this framework, for example, studies in engineering often assess resilience by measuring the cumulative loss in infrastructure quality over time (Fang et al., 2022; Chen et al., 2024; Sahebjamnia et al., 2015). In economics, resilience is often understood as an economic system’s dynamic response to external shocks, and can be thus measured through response functions (Klimek et al., 2019; Brunnermeier, 2024). In ecology, resilience is typically quantified in terms of recovery rate or the time required to adapt to changing disturbance regimes (Seidl et al., 2016). These studies emphasize the system’s speed of recovery and the magnitude of loss, offering valuable insights into resilience measurement.
In addition to performance-based approaches, scholars also focus on how a system’s structural properties shape its resilience, where structure-based approaches are widely used (Hosseini et al., 2016). Indeed, the choice of approach often reflects which principles of resilience are emphasized, yet no consensus emerges. For example, some studies regard diversification as the most fundamental principle of resilience (Hertel et al., 2021) due to their strong correlations (Essuman et al., 2023; Stevens and Teal, 2024; Habibi et al., 2025; Arndt and Helming, 2025), as higher diversification implies that a system is less likely to collapse in the face of disturbances (Nicholson et al., 2021). Redundancy has also been mentioned as a basic principle of resilience, especially in supply chains, where sufficient back-up facilities enhance supply chain resilience during disruptions (Kamalahmadi and Parast, 2016; Tan et al., 2019; Ivanov, 2024a). In addition, the existing literature also highlights system cohesion, clustering, and efficiency as important elements of resilience (Bhamra et al., 2011; Kahiluoto et al., 2020; Yu et al., 2024), which further complement the principles of resilience.
@@
Furthermore, given that most real-world systems are structured as networks with complex topologies (Barabási and Albert, 1999), a substantial body of literature assesses system resilience by analyzing the evolution of a network’s topological metrics, as a key stream of structure-based approaches. For example, resilience can be assessed using the inverse of the harmonic mean of the shortest path lengths between nodes, referred to as global efficiency (which differs from the notion of efficiency used in our study) (Latora and Marchiori, 2001; Osei-Asamoah and Lownes, 2014; Testa et al., 2015; Xie et al., 2021; Wei et al., 2022; Chen et al., 2024). Moreover, Dixit et al. (2020) measure supply chain resilience by integrating four topological indicators, including density, centrality, size, and connectivity, which is also applied to assess resilience in crude oil and gas trade networks (Mou et al., 2020b; Sun et al., 2023). Notably, within a framework proposed by Gao et al. (2016), existing literature can infer whether a system is resilient based on dynamic equations (Xu et al., 2019; Zhao et al., 2024; Tu et al., 2019), which advances structure-based approaches by incorporating system dynamics.
Now, let us concentrate our attention on the field of food systems. Within this, resilience is generally defined as the ability to maintain a stable food supply—in both quantity and diversity—even under disruptions (Tendall et al., 2015), which has attracted considerable scholarly attention (Zurek et al., 2022; Seekell et al., 2017). Analogously, in the field of food trade systems, resilience is generally defined as the ability to maintain a stable availability of, and access to, food supply via international trade—in both quantity and diversity—even under disruptions (Mena et al., 2022; Jafari et al., 2024). With this definition, both performance-based and structure-based approaches have been employed to measure the resilience of food trade systems. For example, under simulated shock scenarios, Fair et al. (2017) and Gutiérrez-Moya et al. (2021) assess the resilience of wheat trade networks, while Gephart et al. (2016) develop a forward shock-propagation model to evaluate the vulnerability of the global seafood trade network. Furthermore, Laber et al. (2023) and Baum et al. (2025) employ multilayer network approaches to examine how shocks propagate through intertwined production and trade channels, while Kuhla et al. (2025) use an agent-based model to reproduce crisis dynamics in agricultural markets. Employing topological indicators, Lin et al. (2014) identify structural vulnerabilities in U.S. food flow networks, and Ji et al. (2024) analyze the vulnerability of global food trade networks. Moreover, Karakoc and Konar (2021) develop a resilience framework based on eigenvalues and nodes’ dependence on major exporters. Similar approaches are also applied in other studies (Wang and Dai, 2021; Grassia et al., 2022; Xu et al., 2024).
Despite extensive academic discussions on the resilience of food trade systems, existing related studies remain incomplete and in need of further complement. First, the principles that constitute resilience remain unsettled, leading to ongoing debates about how it should be measured (Kahiluoto et al., 2020). Although diversification and redundancy are often regarded as the most fundamental principles of resilience, research on supply chains suggests that their single-minded pursuit may erode profits without improving resilience (Pettit et al., 2019). Instead, an emerging consensus among both scholars and practitioners emphasizes that enhancing resilience requires achieving an appropriate trade-off among multiple principles (Brede and de Vries, 2009; Kahiluoto et al., 2020; Yang et al., 2024), a perspective that warrants further empirical substantiation (Kamalahmadi and Parast, 2016).
Second, existing approaches remain inadequate. On the one hand, performance-based approaches emphasize system performance after disturbances while neglecting baseline conditions in the absence of shocks, which obscures the fact that resilience is a fundamental system property (Liu et al., 2022). In addition, determining shock duration and recovery thresholds remains challenging and often unrealistic (Hosseini et al., 2016). On the other hand, structure-based measures such as topology-based measures often overlook network dynamics and fail to capture the temporal evolution of resilience (Qi and Mei, 2024). There are, of course, notable exceptions. For example, Wei et al. (2025) integrate the Resilience Triangle with dynamic equations in a unified framework, offering new methodological insights. Yet, resilience inference within the framework of Gao et al. (2016) relies on the strong assumption that the degrees of interconnected nodes are largely independent (Liu et al., 2024a). Such an assumption contrasts with the high assortativity observed in international trade networks (Mou et al., 2020a), which may limit the applicability of this framework in the context of global food trade. Consequently, there is a pressing need for resilience measures that not only reflect the dynamic evolution of network structures but also capture the fundamental principles underpinning resilience (Li et al., 2024).
Third, existing research needs to strengthen its focus on the internal mechanisms driving resilience and on the roles of peripheral economies in maintaining overall resilience. On the one hand, many studies primarily examine correlations between food trade system resilience and determinants such as topological indicators and population size (Ji et al., 2024; Gutiérrez-Moya et al., 2021), while few investigate the internal mechanisms that drive resilience. On the other hand, core trade economies have been the main focus through various centrality measures (Grassia et al., 2022; Gutiérrez-Moya et al., 2021; Zhang and Zhou, 2023), yet because many such measures emphasize developed economies with numerous partners, they still potentially overlook economies that have peripheral positions in the network but contribute to maintaining the overall resilience. Moreover, link-level disruptions occur more frequently in real-world trade systems, yet existing research predominantly examines the significance of nodes, leaving the importance of trade links relatively less explored (Qian et al., 2017; Guo et al., 2025).
@@ 3. RQ3:
The subsequent sections of this paper are structured in the following manner. The methodological framework is presented in Section 2. Section 3 describes the international food trade network and its resilience. Then, Section 4 and Section 5 present and analyze the main findings. Section 6 discusses the contributions, implications, limitations, and future outlook. Conclusion is finally presented in Section 7.
@@ 2 Methodology
2.1 Topological indicators of the international food trade system
The international food trade system is usually regarded as a network that exhibits small-world (i.e., high clustering coefficients coupled with short average path lengths) and scale-free (i.e., power-law degree distributions) topological indicators in economies’ connectivity (Fagiolo et al., 2009). A food trade network is a weighted digraph $G=\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V}$ represents the vertex set with $N_{\mathcal{V}}$ economies and $\mathcal{E}$ represents the link set with $N_{\mathcal{E}}$ trade relationships. $f_{ij}$ denotes the volume of trade flow that economy $i$ exports to economy $j$. There are several commonly used topological metrics for characterizing network structure, which are briefly described below.
@@ 2.1 Topological indicators of the international food trade system
    Modularity represents the effectiveness of partitioning a network into communities, defined as $Q=\frac{1}{N_{\mathcal{V}}\langle s\rangle}\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\left(f_{ij}-\frac{s_{i}s_{j}}{N_{\mathcal{V}}\langle s\rangle}\right)\delta(c_{i},c_{j})$, where $\delta(c_{i},c_{j})=1$ if economies $i$ and $j$ belong to the same community and $\delta(c_{i},c_{j})=0$ otherwise. The Louvain algorithm, recognized as one of the fastest and best performing algorithms in comparative analysis (Blondel et al., 2008; Lancichinetti and Fortunato, 2009), is employed for community detection in this paper.
@@ 2.2 Efficiency and redundancy of the international food trade system
To address the above methodological gaps, information theory has been introduced into resilience research. In information theory, entropy measures a system’s uncertainty (or diversity). A higher entropy reflects greater unpredictability and diversity, indicating enhanced adaptability to changing conditions and an improved capacity of the system to absorb shocks (Khakifirooz et al., 2025; Reggiani, 2022). Accordingly, we can adopt joint entropy to characterize a system’s overall structural properties (Rutledge et al., 1976; Ulanowicz, 1979), which is also called as “capacity” for system development (Ulanowicz and Norden, 1990). Given that joint entropy is equal to the sum of mutual information and conditional entropy (e.g., $H(X,Y)=I(X,Y)+H(Y|X)+H(X|Y)$), we further introduce two underlying quantities, efficiency (e.g., $I(X,Y)$) and redundancy (e.g., $H(Y|X)+H(X|Y)$), which demonstrate opposing properties of a system.
Efficiency embodies the flow articulation or constraints within networked configurations, which tends to increase due to preferential interactions between economies (Kharrazi et al., 2017). As economies become more interdependent, trade flows tend to become more concentrated and specialized, thereby enhancing overall system efficiency. In information theory, a higher mutual information indicates a increased reduction in uncertainty when some information is known and a stronger statistical inter-dependence between variables. Denoting $s=\sum\limits_{i=1}^{N_{\mathcal{V}}}s_{i}^{\mathrm{out}}=\sum\limits_{j=1}^{N_{\mathcal{V}}}s_{j}^{\mathrm{in}}=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}f_{ij}$, the efficiency $e$ can be expressed as
@@
which is also called as the average mutual information (Rutledge et al., 1976; Ulanowicz and Norden, 1990) and the larger value occurs when trade flows are highly predictable due to the stronger inter-dependence (i.e., $p_{ij}\gg p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}$ for links with larger volume), where
@@
denotes the proportion of the volume $f_{ij}$ of trade flow from economy $i$ to $j$ to the total trade volume $s$, and $p_{i}^{\mathrm{out}}$ and $p_{j}^{\mathrm{in}}$ respectively denote the proportion of export volume from economy $i$ to the others:
@@
and that of import volume from the others to economy $j$:
@@
Redundancy embodies the diversity of food trade pathways, which is critical for the system’s capacity adapting to changing environmental conditions arising from shocks or disturbances (Luo et al., 2024). A greater number of pathways in food trade allows economies to have greater freedom in the trade market and enhances the system’s robustness in coping with risks. Note that our perspective departs from studies that focus primarily on overall uncertainty and instead emphasizes the trade-off between efficiency and redundancy. In this context, redundancy is conceptualized as the residual uncertainty conditional on partial information, which corresponds to the definition of conditional entropy in information theory. A higher conditional entropy indicates a greater availability of alternative pathways in food trade and a consequently higher system redundancy. Accordingly, the redundancy $r$ can be expressed as (Rutledge et al., 1976; Ulanowicz and Norden, 1990)
$\begin{aligned} r&=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{j\mid i}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln p_{i\mid j}\\ &=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}}-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{j}^{\mathrm{in}}}\\ &=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{f_{ij}}{s}\ln\frac{s_{i}^{\mathrm{out}}s_{j}^{\mathrm{in}}}{f_{ij}^{2}}\end{aligned},$ (5)
where $p_{j\mid i}$ denotes the proportion of the export volume to economy $j$ given that the flow is from economy $i$. Similarly, $p_{i\mid j}$ denotes the proportion of the import volume from economy $i$ given that the flow is to economy $j$. The larger value occurs when trade flows can select more pathways due to the higher uncertainty (i.e., $p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}\gg p_{ij}^{2}$ for links with larger volume).